how many yards of materials from a 24-yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces, each 2 1/4 yards long, are removed?
Answer by jojo14344(1513)

If the total length being 24 yards, we subtract the total removed as stated below:
= 3*3-1/2= 10-1/2 yards +
= 5*2-1/4 = 11-1/4 yards
= 21-3/4 yards USED
So the total remaining, 24-(21-3/4)= 2-1/4 yards, oks.
Thank you,
Jojo